# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ Abstract
Unsupervised Domain Adaptation (DA) consists of adapting a model trained on a labeled source domain to perform well on an unlabeled target domain with some data distribution shift. While many methods have been proposed in the literature, fair and realistic evaluation remains an open question, particularly due to methodological difficulties in selecting hyperparameters in the unsupervised setting. With SKADA-Bench, we propose a framework to evaluate DA methods and present a fair evaluation of existing shallow algorithms, including reweighting, mapping, and subspace alignment. Realistic hyperparameter selection is performed with nested cross-validation and various unsupervised model selection scores, on both simulated datasets with controlled shifts and real-world datasets across diverse modalities, such as images, text, biomedical, and tabular data with specific feature extraction. Our benchmark highlights the importance of realistic validation and provides practical guidance for real-life applications, with key insights into the choice and impact of model selection approaches. SKADA-Bench is open-source, reproducible, and can be easily extended with novel DA methods, datasets, and model selection criteria without requiring re-evaluating competitors. SKADA-Bench is available on GitHub at https://github.com/scikit-adaptation/skada-bench.
@@ 1 Introduction
Given some training –or source– data, supervised learning consists in estimating a function that makes good predictions on target data. However, performance often drops when the source distribution used for training differs from the target distribution used for testing. This shift can be due, for instance, to the collection process or non-stationarity in the data, and is ubiquitous in real-life settings. It has been observed in various application fields, including tabular data [16], clinical data [18], or computer vision [15].
@@
Domain adaptation.  Unsupervised Domain Adaptation (DA) addresses this problem by adapting a model trained on a labeled source dataset –or domain– so that it performs well on an unlabeled target domain, assuming some distribution shifts between the two [2, 41, 42]. As illustrated in Figure 1, source and target distributions can exhibit various types of shifts [36]: changes in feature distributions (covariate shift), class proportions (target shift), conditional distributions (conditional shift), or in distributions in particular subspaces (subspace shift). Depending on the type of shift, existing DA methods attempt to align the source distribution closer to the target using reweighting [50, 46], mapping [52, 7], or dimension reduction [39, 12] methods. More recently, it has been proposed to mitigate shifts in a feature space learned by deep learning [15, 53, 29, 9], primarily focusing on computer vision applications. Regardless of the core algorithm used to address the domain shift, hyperparameters must be tuned for optimal performance. Indeed, a critical challenge in applying DA methods to real-world cases is selecting the appropriate method and tuning its hyperparameters, especially given the unknown shift type and the absence of labels in the target domain.
Figure 1: Illustration of the different data shifts studied in the DA literature using the simulated datasets used in the numerical experiments.
@@
Model selection in DA settings.  Without distribution shifts, classical model selection strategies –including hyperparameter optimization– rely on evaluating the generalization error with an independent labeled validation set. However, in DA, validating the hyperparameters in a supervised manner on the target domains is impossible due to the lack of labels. While it is possible to validate the hyperparameters on the source domain, it generally leads to a suboptimal model selection because of the distribution shift. In the literature, this problem is often raised but not always addressed. Some papers choose not to validate the parameters [39], while others validate on the source domain [52] or propose custom cross-validation methods [51]. Few papers focus specifically on DA model selection criteria, which we will call scorers in this paper. These scorers are used to select the methods’ hyperparameters, and mainly consists of reweighting methods on source [49, 59], prediction entropy [37, 45] or circular validation [3]. One of the goals of our benchmark is to evaluate these approaches in realistic scenarios.
@@
Benchmarks of DA.  As machine learning continues to flourish, new methods constantly emerge, making it essential to develop benchmarks that facilitate fair comparisons [22, 31, 35, 40]. In DA and related fields, several benchmarks have been proposed. Numerous papers focus on Out-of-distribution (OOD) datasets for different modalities: computer vision, text, graphs [25, 44], time-series [14], AI-aided drug discovery [23] or tabular dataset [16]. Due to the type of data considered, existing benchmarks are mainly focused on Deep DA methods [38, 56, 24, 11], offering an incomplete evaluation of DA literature. Moreover, only a few benchmarks propose a comparison of Deep unsupervised DA methods with realistic parameters selection, on computer vision [20, 38] and time series [11] data. Those benchmarks have shown the importance of validating with unsupervised scores and reveal that deep DA methods achieve much lower performance in realistic scenarios. In the present work, we focus on “shallow” DA methods, addressing a gap in the DA literature and its evaluation.
@@
Contributions.  In the following, we propose SKADA-Bench, an ambitious and fully reproducible benchmark with the following features: 1. a set of 4 simulated and 8 real-life datasets with different modalities (CV, NLP, tabular, biomedical) totaling 51 realistic shift scenarios, 2. a wide range of 20 shallow DA methods designed to handle different types of shifts, 3. a realistic model selection procedure using 5 different unsupervised scorers with nested cross-validation for hyperparameter selection, 4. an open-source code and publicly available datasets, easy to extend for new DA methods and datasets without the need to re-run the whole experiment.
@@ 2 Domain adaptation and model selection without target labels
In this section, we first discuss the specificities of the unsupervised domain adaptation problem and introduce several types of data shift and their corresponding DA methods. Next, we discuss the different validation strategies used in the literature and the need for realistic scorers to compare DA methods.
@@ 2.1 Data shifts and DA strategies
The theoretical framework of DA is well established [2, 41, 42]. The main results highlight that the performance discrepancy of an estimator between the source and target domains is linked to the divergence between both distributions. This has motivated the majority of DA methods to search for a universal (or domain invariant) predictor by minimizing the divergence between the two domains through the adaptation of the distributions. This is done in practice by modeling and estimating the shift between the source and target distributions and then compensating for this shift before training a predictor.
Data Shifts and DA methods.
@@
A wide variety of shifts between the source and target distributions are possible. They are usually expressed as a relation between the joint distributions $P^{s}(x,y)=P^{s}(x|y)P^{s}_{\mathcal{Y}}(y)=P^{s}(y|x)P^{s}_{\mathcal{X}}(x)$ in the source domain and $P^{t}(x,y)=P^{t}(x|y)P^{t}_{\mathcal{Y}}(y)=P^{t}(y|x)P^{t}_{\mathcal{X}}(x)$ in the target domain. We now discuss the main types of shifts and the strategies proposed in the literature to mitigate them. Figure 1 illustrates these shifts.
@@
In Covariate shift the conditionals probabilities are equal (i.e., $P^{s}(y|x)=P^{t}(y|x)$), but the feature marginals change (i.e., $P^{s}_{\mathcal{X}}(x)\neq P^{t}_{\mathcal{X}}(x)$). Target shift is similar, but the label marginals change $P^{s}_{\mathcal{Y}}(y)\neq P^{t}_{\mathcal{Y}}(y)$ while the conditionals are preserved. For classification problems, it corresponds to a change in the proportion of the classes between the two domains. Both of those shifts can be compensated by reweighting methods that assign different weights to the samples of the source domain to make it closer to the target domain [50, 46].
@@
In Conditional shift, conditional probabilities differ between domain (i.e., $P^{s}(x|y)\neq P^{t}(x|y)$ or $P^{s}(y|x)\neq P^{t}(y|x)$). This shift is typically harder to compensate for, necessitating explicit modeling to address it effectively. For instance, several approaches model the shift as a mapping $m$ between the source and target domain such that $P^{s}(y|m(x))=P^{t}(y|x)$ [52, 7]. The estimated mapping is then applied to the source data before training a predictor.
@@
Subspace shift, also known as domain invariant representation, assumes that while probabilities are different between the domains, there exists a subspace $\mathcal{Z}$ and a function $\phi:\mathcal{X}\to\mathcal{Z}$ such that $P^{s}(y|\phi(x))=P^{t}(y|\phi(x))$. This implies that a classifier trained on this subspace will perform well across both domains. Subspace methods aim to identify the subspace $\mathcal{Z}$ and the function $\phi$, as developed in [39, 12]. Note that, as discussed in the introduction, a natural extension of this idea is to learn a non-linear invariant subspace using deep learning [15, 53].
@@ 2.2 DA model selection strategies
As seen above, the different DA methods are usually designed for one type of shift, yet in a practical problem, one does not know what shift is present. This raises the problem of method and parameter selection when facing a new problem. In this section, we discuss the validation strategies proposed in the literature to compare DA methods, focusing on realistic scorers that do not use target labels.
Realistic DA scorers.
In the literature, few papers propose realistic DA scorers to validate the parameters of the methods, i.e., unsupervised scorers that do not require target labels. The Importance Weighted (IW) scorer [49] computes the score as a reweighted accuracy on labeled sources data. The Deep Embedded Validation (DEV) [59] can be seen as an IW in the latent space with a variance reduction strategy. DEV was originally proposed for deep learning models but can be used on shallow DA methods that compute features from the data (mapping/subspaces). The Prediction Entropy (PE) scorer [37] measures the uncertainty associated with model predictions on the target data. Soft Neighborhood Density (SND) [45] also computes an entropy but on a normalized pairwise similarity matrix between probabilistic predictions on target. The Circular Validation (CircV) scorer [3] performs DA by first adapting the model from the source to the target domain and predicting the target labels. Next, it adapts back from the target to the source using these estimated labels. Performance is measured as the accuracy between the recovered and true source labels.
DA validation in the literature.
The model selection problem in DA has been widely discussed in the literature. Yet, this literature constitutes a subfield of DA and has seldom been used to validate new DA methods. Indeed, there is no consensus on the best validation strategy and many papers do not properly validate their methods, leading to over-estimated performances. Some authors do not discuss the validation procedure [50, 46] or consider fixed hyperparameters [21]. While some methods rely on custom validation technique [51], others use cross-validation, either on the source or the target [52, 7], or alternatively other validation strategies proposed in the literature [6, 3]. A complete picture of the model selection procedures used to validate the methods considered in SKADA-Bench in their original papers is presented in Table A.1 in Appendix. The goal of SKADA-Bench is therefore to constitute a dedicated benchmark to compare scorers from the literature and report performances that can be expected in real use cases for the considered methods.
@@ 3.1 Nested cross-validation loop and implementation
For both loops, the data is split randomly 5 times using stratified sampling with an 80%/20% train/test split. For one given method, we evaluate all the unsupervised scorers discussed earlier, as well as a supervised scorer that uses target labels, over all the nested splits. After averaging, the scores over the splits, the best hyperparameters are selected according to each scorer and then used to train a final classifier on the outer training sets. Although the supervised scorer cannot be used in practice, it is included in our results to actually evaluate the performance drop due to the absence of target labels. To limit complexity and perform a fair comparison of the methods, we set a timeout of 4 hours for performing the nested loop.
Base estimator.
@@
Existing domain adaptation methods typically rely on either a base estimator trained on the adapted data or an iterative estimation process to adapt this estimator to the target data. The choice of the base estimator is crucial, as it significantly impacts the final performance. Before validating the hyperparameters of the DA methods, we determined the best estimator for each dataset using a grid-search on the source data. We tested multiple hyperparameters for Logistic Regression, SVM with RBF kernel, and XGBoost [5], selecting the ones that maximize the average accuracy on the source test sets. Note that for some methods that specifically require an SVM estimator (i.e., JDOT and DASVM), we only validate SVM as the base estimator. We validated the base estimator separately from the DA methods parameters to reduce computational complexity and avoid too complex hyperparameter grids that can compromise the reliability of DA scorers.
Best scorer selection and statistical test.
@@
For all methods, we select the best validation scorer as the one that maximizes the averaged accuracy on the target domains for all real datasets. This provides a reasonable and actionable choice of scorer for each DA method for practitioners. For all methods and datasets, we perform a paired Wilcoxon signed-rank test at the $0.05$ level to detect significant gain or drop in performance with respect to the no DA approach, denoted by “Train Src” in the following. The test is done using the accuracy measures of the DA method with the selected scorer and the Train Src for all shifts and outer splits, ensuring between 10 and 60 values depending on the dataset.
Python implementation.
@@
The benchmark code will be made available on GitHub upon publication of the paper.***Our code is available in supplementary materials. Our benchmark is implemented following the benchopt framework [35], which provides standardized ways of organizing and running benchmarks for ML in Python. This framework facilitates reproducing the benchmark’s results, with tools to install the dependencies, run the methods in parallel, or cache the results to prevent redundant computations. It also makes it easy to extend the benchmark with additional datasets and methods, enabling it to evolve to account for the advances in the field. In the supplementary materials, we provide examples demonstrating how to add DA methods or datasets to the benchmark. Using this framework, we aim to make SKADA-Bench a reference benchmark to evaluate new DA methods in realistic scenarios with valid performance estimations.

### v2

@@ Abstract
Unsupervised Domain Adaptation (DA) consists of adapting a model trained on a labeled source domain to perform well on an unlabeled target domain with some data distribution shift. While many methods have been proposed in the literature, fair and realistic evaluation remains an open question, particularly due to methodological difficulties in selecting hyperparameters in the unsupervised setting. With SKADA-Bench, we propose a framework to evaluate DA methods on diverse modalities, beyond computer vision task that have been largely explored in the literature. We present a complete and fair evaluation of existing shallow algorithms, including reweighting, mapping, and subspace alignment. Realistic hyperparameter selection is performed with nested cross-validation and various unsupervised model selection scores, on both simulated datasets with controlled shifts and real-world datasets across diverse modalities, such as images, text, biomedical, and tabular data. Our benchmark highlights the importance of realistic validation and provides practical guidance for real-life applications, with key insights into the choice and impact of model selection approaches. SKADA-Bench is open-source, reproducible, and can be easily extended with novel DA methods, datasets, and model selection criteria without requiring re-evaluating competitors. SKADA-Bench is available on Github at https://github.com/scikit-adaptation/skada-bench.
@@ 1 Introduction
Given some training –or source– data, supervised learning consists in estimating a function that makes good predictions on target data. However, performance often drops when the source distribution used for training differs from the target distribution used for testing. This shift can be due, for instance, to the collection process or non-stationarity in the data, and is ubiquitous in real-life settings. It has been observed in various application fields, including tabular data (Gardner et al., 2023), clinical data (Harutyunyan et al., 2019), or computer vision (Ganin et al., 2016b).
@@
Domain adaptation.  Unsupervised Domain Adaptation (DA) addresses this problem by adapting a model trained on a labeled source dataset –or domain– so that it performs well on an unlabeled target domain, assuming some distribution shifts between the two (Ben-David et al., 2006; Quinonero-Candela et al., 2008; Redko et al., 2022). As illustrated in Figure 1, source and target distributions can exhibit various types of shifts (Moreno-Torres et al., 2012): changes in feature distributions (covariate shift), class proportions (target shift), conditional distributions (conditional shift), or in distributions in particular subspaces (subspace shift). Depending on the type of shift, existing DA methods attempt to align the source distribution closer to the target using reweighting (Sugiyama & Müller, 2005; Shimodaira, 2000), mapping (Sun et al., 2017; Courty et al., 2017b), or dimension reduction (Pan et al., 2011; Fernando et al., 2013) methods. More recently, it has been proposed to mitigate shifts in a feature space learned by deep learning (Ganin et al., 2016b; Sun & Saenko, 2016; Long et al., 2015a; Damodaran et al., 2018b), primarily focusing on computer vision applications. Regardless of the core algorithm used to address the domain shift, hyperparameters must be tuned for optimal performance. Indeed, a critical challenge in applying DA methods to real-world cases is selecting the appropriate method and tuning its hyperparameters, especially given the unknown shift type and the absence of labels in the target domain.
@@
Model selection in DA settings.  Without distribution shifts, classical model selection strategies –including hyperparameter optimization– rely on evaluating the generalization error with an independent labeled validation set. However, in DA, validating the hyperparameters in a supervised manner on the target domains is impossible due to the lack of labels. While it is possible to validate the hyperparameters on the source domain, it generally leads to a suboptimal model selection because of the distribution shift. In the literature, this problem is often raised but not always addressed. Some papers choose not to validate the parameters (Pan et al., 2011), while others validate on the source domain (Sun et al., 2017) or propose custom cross-validation methods (Sugiyama et al., 2007b). Few papers focus specifically on DA model selection criteria, which we will call scorers in this paper. These scorers are used to select the methods’ hyperparameters, and mainly consists of reweighting methods on source (Sugiyama et al., 2007a; You et al., 2019), prediction entropy (Morerio et al., 2017; Saito et al., 2021) or circular validation (Bruzzone & Marconcini, 2010a). One of the goals of our benchmark is to evaluate these approaches in diverse and realistic scenarios.
Figure 1: Illustration of the different data shifts and assumptions studied in the DA literature and used in the simulated datasets used in the numerical experiments. The colors indicate the classes, with solid colors representing target and transparent colors representing source in the target data plot.
@@
Benchmarks of DA.  As machine learning continues to flourish, new methods constantly emerge, making it essential to develop benchmarks that facilitate fair comparisons (Hutson, 2018; Pineau et al., 2019; Mattson et al., 2020; Moreau et al., 2022). In DA and related fields, several benchmarks have been proposed. Numerous papers focus on Out-of-distribution (OOD) datasets for different modalities: computer vision, text, graphs (Koh et al., 2021; Sagawa et al., 2022), time-series (Gagnon-Audet et al., 2023), AI-aided drug discovery (Ji et al., 2023) or tabular dataset (Gardner et al., 2023). Due to the type of data considered, existing benchmarks are mainly focused on Deep DA methods (Musgrave et al., 2021; Wang, 2018; Jiang et al., 2022; Fawaz et al., 2023), offering an incomplete evaluation of DA literature. Moreover, only a few benchmarks propose a comparison of Deep unsupervised DA methods with realistic parameters selection, on computer vision (Hu et al., 2023; Musgrave et al., 2021) and time series (Fawaz et al., 2023) data. Those benchmarks have shown the importance of validating with unsupervised scores and reveal that Deep DA methods achieve much lower performance in realistic scenarios.
@@
Contributions.  In the following, we propose SKADA-Bench, an ambitious and fully reproducible benchmark with the following features: 1. A set of 4 simulated and 8 real-life datasets with different modalities (computer vision, NLP, tabular, biomedical) totaling 51 realistic shift scenarios, 2. A wide range of 20 Shallow DA methods designed to handle different types of shifts, 3. An evaluation of 3 deep DA methods on 4 real-world datasets from the computer vision and biomedical modalities, 4. A realistic model selection procedure using 5 different unsupervised scorers with nested cross-validation for hyperparameter selection, 5. An open-source implementation and publicly available datasets, easy to extend for new DA methods and datasets without the need to re-run the whole experiment.
@@ 2 Domain adaptation and model selection without target labels
In this section, we first discuss the specificities of the unsupervised domain adaptation problem and introduce several types of data shifts and their corresponding DA methods. Next, we discuss the different validation strategies used in the literature and the need for realistic scorers to compare DA methods.
@@ 2.1 Data shifts and DA strategies
The theoretical framework of DA is well established (Ben-David et al., 2006; Quinonero-Candela et al., 2008; Redko et al., 2022). The main results highlight that the performance discrepancy of an estimator between the source and target domains is linked to the divergence between both distributions. This has motivated the majority of DA methods to search for a universal (or domain invariant) predictor by minimizing the divergence between the two domains through the adaptation of the distributions. This is done in practice by modeling and estimating the shift between the source and target distributions and then compensating for this shift before training a predictor.
@@
Data Shifts and DA methods. A wide variety of shifts between the source and target distributions are possible. They are usually expressed as a relation between the joint distributions $P^{s}(x,y)=P^{s}(x|y)P^{s}_{\mathcal{Y}}(y)=P^{s}(y|x)P^{s}_{\mathcal{X}}(x)$ in the source domain and $P^{t}(x,y)=P^{t}(x|y)P^{t}_{\mathcal{Y}}(y)=P^{t}(y|x)P^{t}_{\mathcal{X}}(x)$ in the target domain. We now discuss the main types of shifts and the strategies proposed in the literature to mitigate them. Figure 1 illustrates these shifts.
@@
In Covariate shift the conditionals probabilities are equal (i.e., $P^{s}(y|x)=P^{t}(y|x)$), but the feature marginals change (i.e., $P^{s}_{\mathcal{X}}(x)\neq P^{t}_{\mathcal{X}}(x)$). Target shift is similar, but the label marginals change $P^{s}_{\mathcal{Y}}(y)\neq P^{t}_{\mathcal{Y}}(y)$ while the conditionals are preserved. For classification problems, it corresponds to a change in the proportion of the classes between the two domains. Both of those shifts can be compensated by reweighting methods that assign different weights to the samples of the source domain to make it closer to the target domain (Sugiyama & Müller, 2005; Shimodaira, 2000).
@@
In Conditional shift, conditional probabilities differ between domain (i.e., $P^{s}(x|y)\neq P^{t}(x|y)$ or $P^{s}(y|x)\neq P^{t}(y|x)$). This shift is typically harder to compensate for, necessitating explicit modeling to address it effectively. For instance, several approaches model the shift as a mapping $m$ between the source and target domain such that $P^{s}(y|m(x))=P^{t}(y|x)$ (Sun et al., 2017; Courty et al., 2017b). The estimated mapping is then applied to the source data before training a predictor.
@@
Subspace shift assumes that while probabilities are different between the domains ($P^{s}_{\mathcal{X}}(x)\neq P^{t}_{\mathcal{X}}(x)$ and $P^{s}(x|y)\neq P^{t}(x|y)$), there exists a subspace $\mathcal{Z}$ and a function $\phi:\mathcal{X}\to\mathcal{Z}$ such that $P^{s}_{\mathcal{Z}}(\phi(x))=P^{t}_{\mathcal{Z}}(\phi(x))$ and $P^{s}(y|\phi(x))=P^{t}(y|\phi(x))$. Note that this means the shift occurs in the orthogonal complement of $\mathcal{Z}$. This implies that a classifier trained on $\mathcal{Z}$ will perform well across both domains. Subspace methods are specifically designed towards identifying the subspace $\mathcal{Z}$ and the function $\phi$, as developed in Pan et al. (2011); Fernando et al. (2013). Note that, as discussed in the introduction, a natural extension of this idea is to learn an invariant feature space using Deep learning (Ganin et al., 2016b; Sun & Saenko, 2016).
@@ 2.2 DA model selection strategies
As seen above, DA methods are typically designed to correct a specific type of shift. However, in real-world scenarios, the nature of the shift is often unknown. This presents a challenge in selecting the appropriate method and tuning its parameters when facing a new problem. In this section, we discuss the validation strategies proposed in the literature to compare DA methods, focusing on realistic scorers that do not use target labels.
Realistic DA scorers. In the literature, few papers propose realistic DA scorers to validate the parameters of the methods, i.e., unsupervised scorers that do not require target labels. The Importance Weighted (IW) scorer (Sugiyama et al., 2007a) computes the score as a reweighted accuracy on labeled sources data. The Deep Embedded Validation (DEV) (You et al., 2019) can be seen as an IW in the latent space with a variance reduction strategy. DEV was originally proposed for Deep learning models but can be used on shallow DA methods that compute features from the data (mapping/subspaces). The Prediction Entropy (PE) scorer (Morerio et al., 2017) measures the uncertainty associated with model predictions on the target data. Soft Neighborhood Density (SND) (Saito et al., 2021) also computes an entropy but on a normalized pairwise similarity matrix between probabilistic predictions on target. The Circular Validation (CircV) scorer (Bruzzone & Marconcini, 2010a) performs DA by first adapting the model from the source to the target domain and predicting the target labels. Next, it adapts back from the target to the source using these estimated labels. Performance is measured as the accuracy between the recovered and true source labels.
The MixVal scorer (Hu et al., 2023) also performs domain adaptation by first adapting the model from the source to the target domain and predicting the target labels. Then, it generates mixed target samples by probing intra-cluster samples to assess neighborhood density and inter-cluster samples to examine classification boundaries. The score is the accuracy between the generated targets labels and their predictions to evaluate the consistency.
DA validation in the literature. The model selection problem in DA has been widely discussed in the literature. Yet, this literature constitutes a subfield of DA and has seldom been used to validate new DA methods. Indeed, there is no consensus on the best validation strategy and many papers do not properly validate their methods, leading to over-estimated performances. Some authors do not discuss the validation procedure (Sugiyama & Müller, 2005; Shimodaira, 2000) or consider fixed hyperparameters (Huang et al., 2006). While some methods rely on custom validation techniques (Sugiyama et al., 2007b), others use cross-validation, either on the source or the target (Sun et al., 2017; Courty et al., 2017b), or alternatively other validation strategies proposed in the literature (Courty et al., 2017a; Bruzzone & Marconcini, 2010a). A complete picture of the model selection procedures used to validate the methods considered in SKADA-Bench in their original papers is presented in Table 4 in Appendix B. The goal of SKADA-Bench is therefore to constitute a dedicated benchmark to compare scorers from the literature and report performances that can be expected in real use cases for the considered methods.
@@ 3.1 Nested cross-validation loop and implementation
For both loops, the data is split randomly 5 times using stratified sampling with an 80%/20% train/test split, except for Deep DA methods, where only one split is computed for the outer loop due to computation time. For one given method, we evaluate all the unsupervised scorers discussed earlier, as well as a supervised scorer that uses target labels, over all the nested splits. After averaging, the scores over the splits, the best hyperparameters are selected according to each scorer and then used to train a final classifier on the outer training sets. Although the supervised scorer cannot be used in practice, it is included in our results to actually evaluate the performance drop due to the absence of target labels. To limit complexity and perform a fair comparison of the methods, we set a timeout of 4 hours for performing the nested loop. Additionally, we chose not to use the CircV scorer for Deep DA methods, as training neural networks twice is computationally expensive.
@@
Base estimators and neural networks. Existing shallow domain adaptation methods typically rely on either a base estimator trained on the adapted data or an iterative estimation process to adapt this estimator to the target data. The choice of the base estimator is crucial, as it significantly impacts the final performance. Before validating the hyperparameters of the DA methods, we determined the best estimator for each dataset using a grid-search on the source data. We tested multiple hyperparameters for Logistic Regression, SVM with RBF kernel, and XGBoost (Chen & Guestrin, 2016), selecting the ones that maximize the average accuracy on the source test sets. Note that for some methods that specifically require an SVM estimator (i.e., JDOT and DASVM), we only validate SVM as the base estimator. We validated the base estimator separately from the DA methods parameters to reduce computational complexity and avoid too complex hyperparameter grids that can compromise the reliability of DA scorers. For Deep DA methods, we similarly select an appropriate architecture and experimental setup for training on the source data for each dataset: a two-layer convolutional neural network for MNIST/USPS, a ResNet50 (He et al., 2016) pretrained on ImageNet (Deng et al., 2009) for Office31 and Office Home, and a ShallowFBCSPNet (Schirrmeister et al., 2017) for BCI. These architectures are widely used and well-supported in the literature of computer vision (Musgrave et al., 2021) and BCI (Schirrmeister et al., 2017). During the nested loop, only the DA parameters for each method are validated.
@@
Best scorer selection and statistical test. For all methods, we select the best validation scorer as the one that maximizes the averaged accuracy on the target domains for all real datasets. This provides a reasonable and actionable choice of scorer for each DA method for practitioners. For all methods and datasets, we perform a paired Wilcoxon signed-rank test at the $0.05$ level to detect significant gain or drop in performance with respect to the no DA approach, denoted by “Train Src” in the following. The test is done using the accuracy measures of the DA method with the selected scorer and the Train Src for all shifts and outer splits, ensuring between 10 and 60 values depending on the dataset. Note that these statistical tests are not performed on Deep DA methods, as the number of splits is too limited for meaningful testing.
@@
Python implementation. The benchmark code is available on GitHub upon publication of the paper.***Our code is available in supplementary materials. Our benchmark is implemented following the benchopt framework (Moreau et al., 2022), which provides standardized ways of organizing and running benchmarks for ML in Python. This framework facilitates reproducing the benchmark’s results, with tools to install the dependencies, run the methods in parallel, or cache the results to prevent redundant computations. It also makes it easy to extend the benchmark with additional datasets and methods, enabling it to evolve to account for the advances in the field. In the supplementary materials, we provide examples demonstrating how to add DA methods or datasets to the benchmark. Using this framework, we aim to make SKADA-Bench a reference benchmark to evaluate new DA methods in realistic scenarios with valid performance estimations.